A plane traveling 135 mph overtakes a speeding car traveling 75 mph, in the same direction. The car had a 2 hour and 45 minute head start. How far from the starting point does the plane overtake the car?
Answer by mananth(16946)

A plane traveling 135 mph overtakes a speeding car traveling 75 mph, in the same direction. The car had a 2 hour and 45 minute head start. How far from the starting point does the plane overtake the car?

car speed = 75 m/h
head start = 2 .75 hours
distance it has already traveled = 2.75*75
=206.25 km
plane speed = 135 mph
car speed = 75 mph
catch up speed = 135-75=60 mph
time taken = 206.75/60
3.45 hours
speed of plane = 135 mph
time = 3.75
distance where it catches up = 3.75*135
506.25 miles